A coffee merchant has coffee beans that sell for $9 per
pound and $12 per pound. The two types are to be mixed to create 100 lb of a mixture
that will sell for $11.25 per pound. How much of each type of bean should be
used in the mixture?

Answer by ptaylor(2198)

Let x= amount of $12 coffee
Then 100-x=amount of $9 coffee
Value of $12 coffee (12x) plus value of $9 coffee (9(100-x)) has to equal value of final mixture $11.25(100). So our equation to solve is:

12x+9(100-x)=11.25(100) get rid of parens
12x+900-9x=1125 subtract 900 from both sides
3x=1125-900
3x=225 divide both sides by x
x=75 lbs----------------------------------amount of $12 coffee
100-x=100-75=25 lbs---------------------------amount of $9 coffee
Ck
75*12+25*9=100*(11.25)
900+225=1125
1125=1125


Hope this helps------ptaylor